simplify the expression : (15x+21)/(9x+15)
*look for an integer factor thats common to numerator and denomenator
seperate the numerator and denomenator into sections: 15x/9x + 21/15
find a common denomenator therefore, 15x/9x * (15/15) and 21/15* (9x/9x); now add together
225x+189x/135x = 364x/135
anyways, i tried this but it did not give me the right solution. any insight?

Answer by John10(297)

You can not simplify like the way that you did. You have to factor the numerator and denominator separately as below:
15x + 21 = 3(5x + 7)
9x + 15 = 3(3x + 5)
To simplify: just cancel out the constant term "3" from the denominator and numerator.
So (15x + 21)/(9x + 15) = (5x + 7)/(3x + 5)
John10:)